# Unlabeled Data or Pre-trained Model: Rethinking Semi-Supervised Learning and Pretrain-Finetuning

### v3

@@ 5 Discussion
Effects of training with long-tailed data on pre-trained models. When training data follows a long-tailed distribution, excessive fine-tuning of pre-trained models can be detrimental, particularly to the performance on tail classes. As highlighted in [47], aggressive fine-tuning can disrupt the class-conditional structure learned during pre-training, resulting in poor representation for underrepresented classes. To address this issue, they proposed LIFT, a lightweight fine-tuning approach that modifies only a small portion of the model (e.g., prompts or adapter layers), effectively adapting to long-tailed distributions while preserving the semantic consistency of the pre-trained model. This demonstrates that leveraging pre-trained models can help mitigate the adverse effects of long-tailed distributions. However, current approaches primarily focus on reducing fine-tuning intensity, often at the cost of head class performance. This highlights the need for strategies that can guide models to fully adapt to imbalanced data while minimizing learning bias.
6 Conclusion and Recommendation
Our experiments provide a comprehensive and fair analysis of the strengths and weaknesses of SSL and pretrain-finetuning paradigms. When should each approach be preferred in real-world applications, and what are their advantages and limitations? Our key insights are as follows:
  • The SSL Paradigm: Best suited for scenarios with low-resolution images or ambiguous class semantics. However, it struggles with noisy, unlabeled data, often requiring more training time and a larger number of training parameters.
  • The Pretrain-Finetuning Paradigm: Ideal for handling imbalanced data distributions and noisy, unlabeled datasets, but tends to underperform on low-resolution images due to its heavy reliance on high-resolution visual inputs and clear semantic information.
Our study also points to promising directions for future research, including but not limited to designing hybrid approaches that effectively combine the strengths of both paradigms:
  • Datasets: Higher-resolution and more challenging datasets should be included to better evaluate the performance of semi-supervised algorithms and pre-trained models, in line with the continued advancement of image resolution.
@@ 6 Conclusion and Recommendation
  • The SSL Paradigm: Future work on semi-supervised learning for image classification should include comparisons with pre-trained models (e.g., VLMs) under matched labeled data conditions.
  • The Pretrain-Finetuning Paradigm: Pre-trained models should be made more robust to low-resolution images to enhance their generalization across a wider range of real-world scenarios.
  • General Directions: Explore methods that integrate semi-supervised learning with pre-trained model adaptation to further improve their performance in labeled data-scarce environments.
One limitation of our study is that we focused solely on evaluating CLIP-based pre-trained models for image classification. In the future, we will conduct more extensive comparisons, hoping that our results and findings will inspire further advancements in SSL and Pretrain-Finetuning paradigms.
Appendix A Detailed Experimental Settings
A.1 Overal Experimental Settings
@@
A.2 Detailed Semi-supervised Datasets and Experimental Setup
Datasets Setup. We evaluate the performance of both classical SSL methods and VLMs on standard image classification benchmarks: CIFAR-10, CIFAR-100 [27], STL-10 [9], and ImageNet [10]. Each dataset is tested under multiple supervision levels, typically using 40, 250, or 4000 labeled samples depending on the benchmark. We also consider different augmentation strategies (e.g., weak/strong) following the standard SSL protocol.
Experimental Setup. For classical SSL baselines (FixMatch [48], MixMatch [5], ReMixMatch [4]), models are pre-trained on the labeled subset and optimized using a combination of supervised and consistency-based unsupervised losses. We directly use results reported in the FixMatch paper to maintain consistency. The unlabeled data is drawn from the remaining training set without labels. For Vision-Language Models (VLMs), we adopt the ViT-B/16 architecture across all experiments. Since VLMs are pre-trained on large-scale image-text pairs, we do not require any additional unlabeled samples. Instead, we fine-tune them using only the labeled portion of the data, treating the task as a pure few-shot classification problem. We use prompt tuning methods such as CoOp and PromptSRC (depending on the experiment), and all models are trained for 50 epochs using the labeled data only. Hyperparameters are kept consistent with those used in the Base-to-New setting. For CoOp, we follow the settings reported in the original paper: 4-shot training for 100 epochs and 25/100/400-shot training for 200 epochs. PromptSRC strictly follows the configurations reported in its paper, with key parameters as follows: both textual and visual prompts consist of 4 tokens, and the first nine transformer layers are modified with learnable prompt tokens. All PromptSRC experiments are run for 20 epochs. Experiments are conducted on a single NVIDIA A800 GPU.
A.3 Detailed Open-set Semi-supervised Datasets and Experimental Setup
@@
  • ImageNet-30. This task uses the ImageNet-30 dataset [22], a 30-class subset of ImageNet. Following [42], we select 5% of the data (approximately 50 images per class) from the first 20 alphabetically sorted classes as labeled data, and treat the rest as unlabeled. For VLMs, the minimum number of samples among these 50 classes is 754; 10% of that equals roughly 70. To maintain consistency and simplify comparisons, we fix the labeled set to 50-shot per class for fine-tuning.
@@
  • ImageNet-100. This task is based on the ImageNet-100 dataset, a 100-class subset of ImageNet as defined in [6]. The classes are split evenly into 50% in-distribution (ID) and 50% out-of-distribution (OOD) based on alphabetical order. For the labeled data, we select 50 images per class from the first 20 ID classes, while the remaining data from all 100 classes is used as unlabeled data. VLMs are fine-tuned using only the 50-shot labeled data from the first 20 ID classes, and evaluated on the corresponding test set following the same setup.
@@
Experimental Setup. Since VLMs do not utilize unlabeled data, the Open-set semi-supervised setting on above datasets essentially becomes a few-shot learning problem with a reduced number of classes. The results of OSSL methods refer to DWD-SL [3].
@@
A.4 Detailed Open-World Semi-Supervised Datasets and Experimental Setup
@@
Dataset Setup. We evaluate our approach on CIFAR-10/100 [27], ImageNet-100 [6], and Tiny ImageNet [31]. Specifically, the ImageNet-100 dataset contains 100 classes sub-sampled from ImageNet-1k following the selection in [10]. For all datasets, we split classes into seen and new categories with a predefined new class ratio (defaulting to 50% unless otherwise specified). From the seen classes, a fixed ratio of data is randomly selected as labeled data, while the remaining seen samples, together with all samples from the new classes, are assigned to the unlabeled set.
@@
Experimental Setup. Under the Base-to-New setting commonly used for evaluating VLMs [25, 57], each dataset’s classes are split evenly into base and new groups. VLMs are trained exclusively on the base classes and then evaluated separately on both base and new classes to assess their generalizability. OWSSL setting closely aligns with the Base-to-New setting, but differs in how unseen (new) classes are treated during training: SSL methods assume the unlabeled pool contains OOD (new) samples and employ clustering or OOD-detection mechanisms to identify and leverage them. While VLMs use only labeled base-class data, without any exposure to OOD examples. At evaluation time, VLMs simply rely on their pre-trained image–text alignment to generalize to new categories, whereas OWSSL methods benefit from the self-label strategy during training [39, 45].
@@
We report OWSSL results using OwMatch [39]. For VLMs, we adopt ViT-B/16 backbones across all variants. CoOp and PromptSRC are fine-tuned using only the labeled data under varying label ratios (e.g., 1%, 5%, 10%), and are trained for 50 epochs following [25, 57]. All other hyperparameters remain consistent across datasets and methods unless otherwise stated.
A.5 Detailed Long-Tailed Datasets and Experimental Settings
@@
Dataset Setup. We evaluate methods on widely-used long-tailed SSL benchmarks: CIFAR-10-LT, CIFAR-100-LT [27], STL-10-LT [9], and ImageNet-127 [12]. For synthetic datasets, the number of samples for each class $c$ is defined as $N_{c}=N_{1}\cdot\gamma_{l}^{-\frac{c-1}{C-1}}$, where $N_{1}$ is the number of samples in the most frequent class and $\gamma_{l}$ is the imbalance ratio. Unlabeled data follows a similar distribution with $\gamma_{u}$.
@@
  • CIFAR-10-LT: Following DASO [52], we adopt $(N_{1},M_{1})\in\{(500,4000),(1500,3000)\}$ and $\gamma_{l}=\gamma_{u}\in\{100,150\}$. We also consider uniform/reversed unlabeled distributions with $\gamma_{u}\in\{1,1/100\}$ and $\gamma_{l}=100$ fixed.
    Based on the minimum labeled class size, we define the corresponding few-shot settings for VLMs:
@@
  • ImageNet-127: This is a real-world long-tailed dataset with naturally imbalanced label distribution. Following CCL [58], we sample 10% of training data for the labeled set. Due to this sampling strategy, there is no fixed $\gamma$ value.
Experimental Setup. Since the LTSSL setting only affects the training set distribution, we evaluate all models on the same balanced test sets as prior work. For example, in CIFAR-10-LT with $\gamma_{l}=\gamma_{u}=100$, $N_{1}=500$ implies that the rarest class has only 5 labeled samples. For VLMs, we treat this as a 5-shot setup and do not use any unlabeled data. All VLMs follow the same architecture (ViT-B/16), hyperparameters, and optimization schedules as in Section 4.2.
Appendix B Detailed CIFAR10/100 Results under Different Backbones
To ensure a consistent amount of training data and other training parameters, we trained PromptSRC with various visual encoders on the CIFAR dataset, running each configuration three times. As shown in Table 7, the experimental results demonstrate a gradual increase in model performance as the visual encoder’s parameter count grows. This suggests that the suboptimal performance on low-resolution images is not attributable to the visual encoder itself.
Table 7: Performance of different visual backbones on the CIFAR-10/100 datasets.
Dataset Backbones
ResNet-50 ViT-B/32 ViT-B/16
CIFAR10 70.23$\pm$0.56 82.00$\pm$0.70 85.10$\pm$0.29
CIFAR100 39.87$\pm$0.48 54.00$\pm$0.29 57.43$\pm$0.25
Appendix C Detailed Analysis and Visualization of ImageNet-127
ImageNet-127 compresses the original 1,000 fine-grained ImageNet classes into 127 higher-level categories, inevitably producing a long-tailed distribution. For example, 124 distinct classes—including abacus, acoustic guitar, and analog clock—are merged into the super-category device, whereas a single class such as bubble constitutes the entire globule category. Thus, categories that absorb many fine-grained classes become head classes, while those represented by only one class form the tail.

### v4

@@ 5 Conclusions and Discussions
Discussion 2: Effects of training with long-tailed data on pre-trained models. When training data follows a long-tailed distribution, excessive fine-tuning of pre-trained models can be harmful, particularly degrading performance on tail classes. Aggressive fine-tuning can disrupt the pre-trained class-conditional structure and weaken rare category representations (Shi et al., 2024). To mitigate this, LIFT fine-tunes only a small part of the model, enabling adaptation to long-tailed distributions while avoiding catastrophic forgetting. However, most existing approaches focus on reducing fine-tuning intensity, often at the cost of head class performance. This highlights the urgent need for new strategies that can adapt to imbalanced data while minimizing learning bias.
6 Recommendations for Future Research
Facing label scarcity, choosing between SSL methods and pre-trained models is a critical decision. Based on our practical considerations, here are our recommendations:
One should first attempt to address the problem using pre-trained models along with fine-tuning strategies. If “adaptation bias” arises, it is advisable to incorporate additional pseudo-labeling to facilitate adaptation. If these approaches prove ineffective and the discrepancy between training and testing classes is relatively small, then it may be beneficial to design a task-specific SSL method—either a traditional SSL approach or one that leverages pre-trained models for pseudo-label generation—to train a specialized model for the downstream task. Our study also highlights five promising directions for future research:
  • Datasets: Incorporate higher-resolution and more challenging benchmarks to better assess SSL methods and pre-trained models, aligning with the steady rise in image resolution.
@@ 6 Recommendations for Future Research
  • SSL methods: Future work on image-classification SSL should include head-to-head comparisons with strong pre-trained baselines under matched conditions (e.g., identical data, resolution, augmentation, and compute).
  • Pre-trained models: Current pre-trained models remain sensitive to distribution shift (e.g., resolution mismatches between pre-training and testing data). More research is needed on robust adaptation strategies and evaluation protocols to mitigate these effects.
  • MLLMs Although MLLMs demonstrate strong semantic understanding capabilities, their limitations in visual encoders lead to challenges in recognizing low-resolution images, which remains an important direction for future research.
  • General directions: Explore principled integrations of SSL with adaption of pre-trained models to enhance performance in label-scarce settings.
One limitation of our study is that, since most SSL methods are primarily designed for image classification tasks, the range of experimental settings we compared is relatively narrow. In future work, we plan to explore the integration of these two approaches (e.g., leveraging the zero-shot classification capability of VLMs for pseudo-labeling and iterative training) and extend our research to broader multimodal scenarios.
Ethics Statement
This work aims to investigate the performance of existing SSL methods and pretrained models. It does not involve human subjects, sensitive personal data. All datasets used are publicly available, and we adhere to the ICLR Code of Ethics.
Reproducibility Statement
We have taken extensive measures to ensure the reproducibility of our results. An anonymous code repository is available at https://anonymous.4open.science/r/Rethinking-SSL-and-Pretrain-Finetuning-5566/, containing implementations of code and running scripts. Details of experimental settings are described in Section 4 and Appendix B. All datasets used are publicly available.
Appendix A Use of Large Language Models (LLMs)
In this work, LLMs were used solely for language polishing and translation; they did not contribute to research ideation, experiment design, or result interpretation. Multimodal LLMs (e.g., Qwen, GPT-4o-min) were employed only as baselines in our experiments. The authors take full responsibility for all content and results.
Appendix B Detailed Experimental Setup
B.1 Overal Experimental Setup
@@
B.2 Detailed Semi-supervised Datasets and Experimental Setup
Datasets Setup. We evaluate the performance of both classical SSL methods and VLMs on standard image classification benchmarks: CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), STL-10 (Coates et al., 2011), and ImageNet (Deng et al., 2009). Each dataset is tested under multiple supervision levels, typically using 40, 250, or 4000 labeled samples depending on the benchmark. We also consider different augmentation strategies (e.g., weak/strong) following the standard SSL protocol.
Experimental Setup. For classical SSL baselines (FixMatch (Sohn et al., 2020), MixMatch (Berthelot et al., 2019b), ReMixMatch (Berthelot et al., 2019a)), models are pre-trained on the labeled subset and optimized using a combination of supervised and consistency-based unsupervised losses. We directly use results reported in the FixMatch paper to maintain consistency. The unlabeled data is drawn from the remaining training set without labels. For Vision-Language Models (VLMs), we adopt the ViT-B/16 architecture across all experiments. Since VLMs are pre-trained on large-scale image-text pairs, we do not require any additional unlabeled samples. Instead, we fine-tune them using only the labeled portion of the data, treating the task as a pure few-shot classification problem. We use prompt tuning methods such as CoOp and PromptSRC (depending on the experiment), and all models are trained for 50 epochs using the labeled data only. Hyperparameters are kept consistent with those used in the Base-to-New setting. For CoOp, we follow the settings reported in the original paper: 4-shot training for 100 epochs and 25/100/400-shot training for 200 epochs. PromptSRC strictly follows the configurations reported in its paper, with key parameters as follows: both textual and visual prompts consist of 4 tokens, and the first nine transformer layers are modified with learnable prompt tokens. All PromptSRC experiments are run for 20 epochs. Experiments are conducted on a single NVIDIA A800 GPU.
B.3 Detailed Open-set Semi-supervised Datasets and Experimental Setup
@@
  • ImageNet-30. This task uses the ImageNet-30 dataset (Hendrycks et al., 2019), a 30-class subset of ImageNet. Following Pham et al. (2021), we select 5% of the data (approximately 50 images per class) from the first 20 alphabetically sorted classes as labeled data, and treat the rest as unlabeled. For VLMs, the minimum number of samples among these 50 classes is 754; 10% of that equals roughly 70. To maintain consistency and simplify comparisons, we fix the labeled set to 50-shot per class for fine-tuning.
@@
  • ImageNet-100. This task is based on the ImageNet-100 dataset, a 100-class subset of ImageNet as defined in Cao et al. (2021). The classes are split evenly into 50% in-distribution (ID) and 50% out-of-distribution (OOD) based on alphabetical order. For the labeled data, we select 50 images per class from the first 20 ID classes, while the remaining data from all 100 classes is used as unlabeled data. VLMs are fine-tuned using only the 50-shot labeled data from the first 20 ID classes, and evaluated on the corresponding test set following the same setup.
@@
Experimental Setup. Since VLMs do not utilize unlabeled data, the Open-set semi-supervised setting on above datasets essentially becomes a few-shot learning problem with a reduced number of classes. The results of OSSL methods refer to DWD-SL (Ban et al., 2024).
@@
B.4 Detailed Open-World Semi-Supervised Datasets and Experimental Setup
@@
Dataset Setup. We evaluate our approach on CIFAR-10/100 (Krizhevsky et al., 2009), ImageNet-100 (Cao et al., 2021), and Tiny ImageNet (Le & Yang, 2015). Specifically, the ImageNet-100 dataset contains 100 classes sub-sampled from ImageNet-1k following the selection in Deng et al. (2009). For all datasets, we split classes into seen and new categories with a predefined new class ratio (defaulting to 50% unless otherwise specified). From the seen classes, a fixed ratio of data is randomly selected as labeled data, while the remaining seen samples, together with all samples from the new classes, are assigned to the unlabeled set.
@@
Experimental Setup. Under the Base-to-New setting commonly used for evaluating VLMs (Khattak et al., 2023b; Zhou et al., 2022b), each dataset’s classes are split evenly into base and new groups. VLMs are trained exclusively on the base classes and then evaluated separately on both base and new classes to assess their generalizability. OWSSL setting closely aligns with the Base-to-New setting, but differs in how unseen (new) classes are treated during training: SSL methods assume the unlabeled pool contains OOD (new) samples and employ clustering or OOD-detection mechanisms to identify and leverage them. While VLMs use only labeled base-class data, without any exposure to OOD examples. At evaluation time, VLMs simply rely on their pre-trained image–text alignment to generalize to new categories, whereas OWSSL methods benefit from the self-label strategy during training (Niu et al., 2024; Rizve et al., 2022).
@@
We report OWSSL results using OwMatch (Niu et al., 2024). For VLMs, we adopt ViT-B/16 backbones across all variants. CoOp and PromptSRC are fine-tuned using only the labeled data under varying label ratios (e.g., 1%, 5%, 10%), and are trained for 50 epochs following Khattak et al. (2023b); Zhou et al. (2022b). All other hyperparameters remain consistent across datasets and methods unless otherwise stated.
B.5 Detailed Long-Tailed Datasets and Experimental Setup
@@
Dataset Setup. We evaluate methods on widely-used long-tailed SSL benchmarks: CIFAR-10-LT, CIFAR-100-LT (Krizhevsky et al., 2009), STL-10-LT (Coates et al., 2011), and ImageNet-127 (Fan et al., 2022). For synthetic datasets, the number of samples for each class $c$ is defined as $N_{c}=N_{1}\cdot\gamma_{l}^{-(c-1)/(C-1)}$, where $N_{1}$ is the number of samples in the most frequent class and $\gamma_{l}$ is the imbalance ratio. Unlabeled data follows a similar distribution with $\gamma_{u}$.
@@
  • CIFAR-10-LT: Following DASO (Wei & Gan, 2023), we adopt $(N_{1},M_{1})\in\{(500,4000),(1500,3000)\}$ and $\gamma_{l}=\gamma_{u}\in\{100,150\}$. We also consider uniform/reversed unlabeled distributions with $\gamma_{u}\in\{1,1/100\}$ and $\gamma_{l}=100$ fixed. Based on the minimum labeled class size, we define the corresponding few-shot settings for VLMs:
@@
  • ImageNet-127: This is a real-world long-tailed dataset with naturally imbalanced label distribution. Following CCL (Zhou et al., 2024), we sample 10% of training data for the labeled set. Due to this sampling strategy, there is no fixed $\gamma$ value.
Experimental Setup. Since the LTSSL setting only affects the training set distribution, we evaluate all models on the same balanced test sets as prior work. For example, in CIFAR-10-LT with $\gamma_{l}=\gamma_{u}=100$, $N_{1}=500$ implies that the rarest class has only 5 labeled samples. For VLMs, we treat this as a 5-shot setup and do not use any unlabeled data. All VLMs follow the same architecture (ViT-B/16), hyperparameters, and optimization schedules as in the Section about SSL.
Appendix C SSL vs. VLM Finetuning in Long-Tailed Semi-Supervised Setup
Method STL10-LT STL10-LT ImageNet-127
$\gamma=10$ $\gamma=20$ $\gamma=\text{None}$
$N_{1}=150$
$M=100k$
$N_{1}=450$
$M=100k$
$N_{1}=150$
$M=100k$
$N_{1}=450$
$M=100k$
32 $\times$ 32 64 $\times$ 64
w/ DARP (Kim et al., 2020) 66.9 $\pm$ 1.66 75.4 $\pm$ 1.05 59.9 $\pm$ 2.17 72.3 $\pm$ 0.60 30.5 42.5
w/ CReST+ (Wei et al., 2021) 56.0 $\pm$ 3.13 71.5 $\pm$ 0.96 55.4 $\pm$ 3.03 68.5 $\pm$ 1.15 32.5 44.7
w/ DASO (Oh et al., 2022) 61.3 $\pm$ 1.84 78.4 $\pm$ 0.96 68.5 $\pm$ 1.78 75.3 $\pm$ 0.44 40.9 55.9
w/ ACR (Wei & Gan, 2023) 70.2 $\pm$ 3.04 83.0 $\pm$ 1.40 75.1 $\pm$ 1.81 80.5 $\pm$ 0.25 57.2 63.6
w/ CPE (Ma et al., 2024) 69.6 $\pm$ 0.46 83.0 $\pm$ 1.14 69.6 $\pm$ 0.46 81.7 $\pm$ 0.34 - -
w/ CCL (Zhou et al., 2024) 79.1 $\pm$ 0.43 84.8 $\pm$ 1.05 77.1 $\pm$ 0.33 83.1 $\pm$ 0.18 61.5 67.8
ZSCLIP (Radford et al., 2021) 97.4 $\pm$ 0.00 46.00 $\pm$ 0.00
CoOp (Zhou et al., 2022b) 97.77 $\pm$ 0.19 98.13 $\pm$ 0.17 97.73 $\pm$ 0.25 97.83 $\pm$ 0.09 49.83 $\pm$ 1.04 55.23 $\pm$ 0.50
PromptSRC (Khattak et al., 2023b) 98.27 $\pm$ 0.05 98.47 $\pm$ 0.05 98.13 $\pm$ 0.12 98.30 $\pm$ 0.08 60.33 $\pm$ 0.17 64.37 $\pm$ 0.34
Experiment Results. As shown in Table 8, pre-trained models perform well on highly imbalanced datasets like STL-10 but struggle on semantically coarse datasets such as ImageNet-127 (Fan et al., 2022). ImageNet-127 is obtained by collapsing the 1,000 fine-grained ImageNet-1K classes into 127 coarse-grained categories according to their positions in the WordNet semantic hierarchy, yielding a long-tailed distribution (e.g., grouping “abacus,” “acoustic guitar,” and “analog clock” under the broader category “device”). Zero-shot CLIP evaluations reveal that this hierarchical coarsening injects semantic ambiguity, making it challenging for pretrained VLMs to confidently associate a fine-grained concept with a specific image. Although downstream fine-tuning can partially mitigate this issue, the underlying semantic constraints remain. By contrast, SSL learns directly from neural representations rather than predefined semantics, making it less vulnerable to such ambiguity and better suited to tasks with limited or ambiguous class structure.
Detailed Analysis and Visualization of ImageNet-127. ImageNet-127 compresses the original 1,000 fine-grained ImageNet classes into 127 higher-level categories, inevitably producing a long-tailed distribution. For example, 124 distinct classes—including abacus, acoustic guitar, and analog clock—are merged into the super-category device, whereas a single class such as bubble constitutes the entire globule category. Thus, categories that absorb many fine-grained classes become head classes, while those represented by only one class form the tail.